helium is found to diffuse two times faster than x, what is the molecular mass of x?
1
Expert's answer
2017-02-21T08:28:40-0500
Grahams law, says:
v1/v2 = sqrt(M2/M1) where v1 and v2 are veloicityes;
M1 and M2 are molecular weights
So is difuse faster in two times
v1/v2 = 2
M(He) = 4 g/mol
2 = sqrt(x/4)
4 = x/4
x = 16 (g/mol)

Answer: 16 (g/mol)